Explain bragg equation of diffraction.


1
Expert's answer
2021-02-15T12:14:37-0500

Bragg's equation of diffraction is expressed as:

"n\lambda =2dsin\theta"

The equation illustrates the reason why the faces of crystals reflect X-ray beams at specific incident angles "\left(\theta ,\lambda \right)".

The variable "d" points the distance between atomic layers while the variable "\lambda" specifically explains the wavelength of incident X-ray beam.

Lastly, "n" is an integer such as: "1,2,3......"